OpenStudy (anonymous):

Which of the following is a possible equivalence that can be used in a conversion? ten cubed kL over one L ten squared daL over one L ten to the negative second power L over one cL ten to the negative first power L over one mL

OpenStudy (anonymous):

The answer is the third one. :) Firstly, when you convert to a different unit, the value - what the number represents - doesn't change. You're actually multiplying the number by 1. Its a fraction, where the top and bottom are equal. I'll derive a simple conversion that I hope you agree with. Even if you're in the stubborn U.S., it's important to know the metric system. Like, 1 meter is one hundred centimters. In math, 1 [m]=100 [cm] This would imply that, by dividing by 1 meter, 1 [m]1 [m]=100 [cm]1 [m] ⇓ 1 [m]1 [m]=100 [cm]1 [m] ⇓ 1=100 [cm]1 [m] That sort of thing is always good to know for unit conversions. And reversing that on your multiple choices will check to see if they are good! 103 [kL][L]→103 [kL]=?[L] 102 [daL][L]→102 [daL]=?[L] 10−2 [L][cL]→10−2 [L]=?[cL] 10−1 [L][mL]→10−1 [L]=?[mL] And note that 10−1=1101 and that 10−2=1102, just by the meaning of negative exponents. Just for my own fun... 103 [kL][L]→103 [kL]=[L] 102 [daL][L]→102 [daL]=[L] 10−2 [L][cL]→10−2 [L]=[cL]✓ 10−1 [L][mL]→10−1 [L]=[mL]